Assume that the proportion of voters who prefer Candidate A is
p
=
0.571
. Organization D conducts a poll of
n
=
5
voters.
Let X represent the number of voters polled who prefer Candidate A. Use some form of appropriate technology (e.g., your calculator or statistics software like Excel, R, or StatDisk) to find the cumulative probability distribution.

Found 2 solutions by ewatrrr, math_tutor2020:
Answer by ewatrrr(24785)


Hi  
Using Excel:
Entered P(x) in spreadsheet
X       P(x)
0	0.0145
1	0.0967
2	0.2574
3	0.3426
4	0.228
5	0.0607
Bar Graph
     
Wish You the Best in your Studies.

Answer by math_tutor2020(3835)


I recommend using a spreadsheet. I'm using LibreOffice but Excel will work the same way.

This is what the answer looks like
kP(X ≤ k)
00.014531
10.111233
20.368654
30.711282
40.939301
51

Take note how the values in the right hand column are always increasing. This is because the CDF values are built upon one another.
Something like P(X ≤ 3) involves P(X ≤ 2)
P(X ≤ 2) = P(X = 1) + P(X = 2)
P(X ≤ 3) = P(X = 1) + P(X = 2) + P(X = 3)
P(X ≤ 3) = P(X ≤ 2) + P(X = 3)
Similar situations happen with the other k values.

Here's what I did to generate that table
In cells A1 and B1 are k and P(X ≤ k) respectively.
In cells A2,A3,...,A7 are 0 through 5

Type this command into cell B2
=BINOMDIST(A2,5,0.571,1)
The first input is the k value from the table.
The second and third inputs are n = 5 and p = 0.571 respectively
The final input 1 tells the spreadsheet to compute the CDF instead of the PDF.
Refer to the documentation for more information.

That above command in blue will compute the CDF value for X = 0
Then double-click the small square in the bottom right corner of that cell. Doing so will auto-fill the remaining CDF values for X = 1, X = 2, up to X = 5.
You should get the table you see above. The rounding precision will vary. I rounded each value to 6 decimal places.

I used the function called ROUND to get that job done.
Examples
=ROUND(1.2345,2) results in 1.23
=ROUND(1.2345,3) results in 1.235
The first input is the value you want to round. The second input is the number of decimal places.

If you want to round and compute the CDF value in one step, then type this in
=ROUND(BINOMDIST(A2,5,0.571,1),6)
Change the '6' at the end to whatever level of rounding precision you want.